(1 point) Suppose that you have purchased a filling machine for candy bags that is supposed to fill each bag with 14 oz of candy. Assume that the weights of filled bags are approximately normally distributed. A random sample of 9 bags yields the following data (in oz): 13.13 13.29 13 13.42 14.52 13.59 14.79 14.59 14.59 Using a 0.05, can you conclude that the mean fill weight is actually less than 14 by computing the following: (a) the degree of freedom (b) the critical t value (c) the test statistics The final conclusion is 0 A. We can reject the null hypothesis that ?-14. 0 B. There is not sufficient evidence to reject the null hypothesis that ?-14. Our results do not provide enough evidence that mean weight is less 14.
Answer #1

a)
df = n - 1 = 9 -1 = 8

b)
critical value = -1.860

c)
t = ( x - mean)/(s/sqrt(n))
= ( 13.88 - 14)/(-/7269/sqrt(9))
= -0.4953

option B)